OpenStudy (anonymous):

If h(x) = x − 1 and j(x) = −4x, find h[j(5)].

OpenStudy (anonymous):

i have more after this can i get your help on more?

OpenStudy (anonymous):

ok well imput a number for x if x equals 1 then h(1)=1-1 and j(1)=-4*1 then hj(5) would be ...... 0*-4(5)

OpenStudy (anonymous):

this means you need to put result of J(5) in h j:5=-4*5=-20 put -20 in h -20-1=-21

OpenStudy (anonymous):

ok ignore the last line of my post

OpenStudy (anonymous):

GUYS I HAVE MORE OKAY"

OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = −x − 2, find (f + g)(−1).